If the sides of a square are decreased by 3cm, the area is decreased by 81 cm^2. What were the dimentions of the original square?
My answer for this one was 15cm by 15cm. Is this correct?
Thanks.

Answer by jim_thompson5910(35256)

Yeah you got this right. You can check by finding the area of a square of length 15 cm. 15*15=225 and find the length of a square of 12 cm: 12*12=144 which is 81 away from 225 (ie 225-144=81)